# Personhood credentials: Artificial intelligence and the value of

### v1

@@
Anonymity is an important principle online. However, malicious actors have long used misleading identities to conduct fraud, spread disinformation, and carry out other deceptive schemes. With the advent of increasingly capable AI, bad actors can amplify the potential scale and effectiveness of their operations, intensifying the challenge of balancing anonymity and trustworthiness online. In this paper, we analyze the value of a new tool to address this challenge: “personhood credentials” (PHCs), digital credentials that empower users to demonstrate that they are real people—not AIs—to online services, without disclosing any personal information. Such credentials can be issued by a range of trusted institutions—governments or otherwise. A PHC system, according to our definition, could be local or global, and does not need to be biometrics-based. Two trends in AI contribute to the urgency of the challenge: AI’s increasing indistinguishability (i.e., lifelike content and avatars, agentic activity) from people online, and AI’s increasing scalability (i.e., cost-effectiveness, accessibility). Drawing on a long history of research into anonymous credentials and “proof-of-personhood” systems, personhood credentials give people a way to signal their trustworthiness on online platforms, and offer service providers new tools for reducing misuse by bad actors. In contrast, existing countermeasures to automated deception—such as CAPTCHAs—are inadequate against sophisticated AI, while stringent identity verification solutions are insufficiently private for many use-cases. After surveying the benefits of personhood credentials, we also examine deployment risks and design challenges. We conclude with actionable next steps for policymakers, technologists, and standards bodies to consider in consultation with the public.
@@
With access to increasingly capable AI, malicious actors can potentially orchestrate more effective deceptive schemes. Two trends contribute to the potential of deceptive schemes:
@@
PHCs improve on and complement existing approaches to countering AI-powered deception online. For example, the following approaches are often not robust to highly capable AI, insufficiently inclusive, and/or not privacy-preserving:
@@ 1.1   The Internet has long struggled with deceptive activity
Defenders of online freedom of expression sometimes view deception as an unfortunate but necessary cost of preserving users’ privacy [32]. Indeed, the Internet’s pioneers saw anonymity as a fundamental pillar of privacy and freedom of expression [148]. They built its architecture to let people participate in digital spaces without disclosing their real identities. The benefits from this approach are numerous, and worthy of steadfast protection—for example, anonymity allows people in oppressive regimes to express their opinions without fear of retribution. However, these benefits came at some cost—one such cost was a lack of accountability for deceptive misuse. [^4] For an account of how these norms evolved on Usenet, see [278]. A common approach to undesired content sharing on Usenet was to apply social pressure [21].
@@
The resulting escalation in deceptive activity across social media, public comment systems, and other essential digital services could create substantial challenges for institutions that rely on the Internet. Under pressure, these institutions may resort to invasive measures for verifying users’ identities online, overturning the Internet’s longstanding commitment to privacy and civil liberties. [^9] One example proposal for associating online activity with individual identity comes from China’s Ministry of Public Security and its Cyberspace Administration [251]. Such a proposal might reduce the amount of individuals’ information available to private companies, but may entail greater information being available to the government.
@@
In this paper, we focus on one potential solution to counter AI-powered deception: personhood credentials (hereafter referred to as “PHCs”), which certify that their holder is a person [^10] For the purposes of this paper, we use “person” to refer to a human being, though we recognize that some domains use the term “person” more broadly. For example, in law, “a person is any being whom the law regards as capable of rights or duties” [223], which can include “artificial persons” (e.g., corporations) in addition to “natural persons” (i.e., human beings) [107]. In moral philosophy, “personhood” need not apply only to human beings and could include other groups like non-human animals; “personhood” is sometimes associated with having full moral status, though philosophers disagree considerably about which traits are necessary or sufficient for full moral status [59]. Though we use “person” and “personhood” to refer to human beings, this should not be read as implying that AI systems should never be considered moral persons or granted legal personhood (whether for moral or pragmatic reasons). without revealing anything more about their identity. PHCs help to distinguish people from even the most advanced AI systems by relying on two important deficits of AI. Specifically, AI systems cannot convincingly mimic people offline, and they cannot bypass state-of-the-art cryptographic systems. [^11] Cryptography relies on computationally hard mathematical problems, such as the factoring of very large numbers. There are not any known methods of efficiently solving certain such problems, whether by a human or an AI system. PHCs allow people to interact with different digital service providers in a way that maintains their privacy through unlinkable pseudonymity—the credential holder can have a persistent, privacy-preserving pseudonym with different service providers, and different service providers have no way to link their activity. The issuer of the credential cannot trace their activity either. [^12] While some national digital identity systems also draw on cryptography, they have very different goals and privacy affordances compared to PHCs. For example, Estonia’s national digital identity card allows citizens to conduct a wide range of activities online (e.g., cryptographically signing important contracts). But these systems are specifically not intended to conceal which holder of a card is performing the action [213].
@@
That personhood credentials ought to exist is not a wholly novel idea. Related credentials are the subject of ongoing initiatives, [^13] Our advocacy of personhood credentials aligns with many ongoing initiatives that aim for minimal disclosure identity and authentication systems, e.g., World Wide Web Consortium’s (W3C’s) Verifiable Credentials [240] and decentralized identifiers [239], European Union Digital Identity’s (EUDI’s) privacy-preserving digital wallets [88], and a range of standards and implementations of anonymous and attribute-based credentials [1, 126, 90]. One particularly relevant implementation is British Columbia’s Person Credential [34, 165]. as well as decades of research in security and cryptography. In particular, a personhood credential is a type of anonymous credential, first proposed in the 1980s, [^14] An anonymous credential is a credential that allows its holder to prove some specific statement about themselves without revealing anything further [45, 47, 158]. In the case of a personhood credential, the claim is that its holder is a person. In other cases, one could have an anonymous credential that proves their age without revealing anything further. Personhood credentials largely coincide with “accountable pseudonyms,” which aim to link a pseudonymous digital certificate that can be used anonymously online to an offline entity [99]. For a recent systematization of knowledge paper about anonymous credentials, see [118]. and can be understood as similar to a “proof-of-personhood” system. [^15] The term “proof of personhood” emerged in blockchain communities—in contrast to the concepts of “proof of work” and “proof of stake”—as a strategy for dealing with Sybil attacks (where one entity subverts a digital system by using many pseudonymous identities to exert undue influence) in permissionless voting systems [29]. Given that “proof of personhood systems” often have specific aims (for instance, to allocate a cryptocurrency or voting rights in a decentralized autonomous organization) and different resulting design goals (e.g., they often aim for global uniqueness, sometimes via biometrics), we use the more generic term “personhood credentials” throughout the paper. What is unprecedented is the scale and urgency of the problems these credentials could address, and the harms to trustworthy interactions [^16] By “trustworthy digital interactions,” we mean interactions in which the parties’ reasonable expectations are achieved: for instance, a trustworthy online ecosystem is one in which service providers are confident that their services are being used as intended, and users can access digital services free from fears of abuse, attack, and other harms. We further unpack the foundations of this concept in A What do we mean by “trustworthy” digital interaction?, drawing largely on [133]. on the Internet that may occur if unmitigated.
@@ 1.2   Trends in AI threaten to make online deception more effective
  • Taking human-like actions across the Internet (e.g., browsing a website, solving CAPTCHAs when challenged, making sophisticated plans to achieve goals they are given).
@@
AI systems are improving in each of these categories. For instance, AI-generated text and images are increasingly difficult to distinguish from human-created content in domains like politics, [^19] For a study of legislators’ response rate to emails drafted by AI compared to those drafted by people—finding only a small difference—see [150]. art [220], and literature [58]. Already, many people struggle to discern whether they are conversing by text with a human or an AI system in some settings. [^20] Research suggests that text written by AI systems are hard to distinguish from text written by people, such as in Turing Test conversations [135, 5]. These findings are not wholly new; at 2020’s GPT-3 launch, OpenAI reported that participants in a research study had trouble distinguishing between news articles by people versus those generated by AI [36]. These abilities may be useful for executing deceptive schemes at scale, as a growing body of literature explores in the context of roleplaying as certain personas [265], making persuasive arguments [201, 17, 14, 85], and tailoring persuasion to a specific recipient [162].
@@
Likewise, AI is improving at creating human-like avatars. AI can enable malicious actors to create realistic high-quality photos and videos of people who do not exist [147, 33, 259, 35]; animate photos of people into videos [233]; take on the appearance of a different person over video chat [276, 123]; and speak with a realistic human voice, whether a generic voice or that of a specific person. [^21] For an overview of voice generation abilities and resultant challenges, see [196]. There are many different types of risks that can emerge with voice cloning [125]. Voice cloning scams, in particular, are rampant and on the rise [4]. These capabilities can be used to spoof identification checks, such as those requiring a selfie holding a matching driver’s license. [^22] Some platform verification processes request proof of a government ID card—but even this process cannot restrict entry of AIs working on behalf of bad actors, as AI tools can be used to spoof a realistic-looking selfie of any person holding their supposed driver’s license [67]. If a service provider is willing to pay for more expensive verifications—such as those offered through the American Association of Motor Vehicle Administrators (AAMVA) in the US—they may be able to filter out some fraudulent IDs, but even these services have limits, such as not offering comparisons against the photo-of-record [7]. In the near future, during a video chat, a user may not be able to discern whether their conversational partner is actually the person they are seeing, a person disguising themselves using AI, or even a complete AI simulation of a real or fictitious person. [^23] There have already been incidents involving convincing deepfaked video calls of executives [50] and politicians [194].
@@ 1.3   Current solutions for countering AI-powered deception need improvement
A second approach is to impose economic barriers that make it costly to perpetuate AI-powered deception at scale. Some websites introduce paid subscriptions with the explicit intent of distinguishing between people and bots by making it more expensive to use many bot accounts. [^35] Elon Musk reported this as one factor in his decision to institute a paid subscription option on Twitter (now X) [224]. These charges, however, disproportionately affect lower-income users and are an insufficient barrier against particularly profitable forms of malicious activity [110, 245]. Relatedly, some digital services check whether users hold a valid credit card, though this approach can be circumvented and also has similar challenges with inclusivity. [^36] Credit card requirements can be circumvented via virtual cards [207].
@@
A fourth approach is to use document- or appearance-based verification to confirm that there is a person behind some digital activity. Some verification protocols require that users join a video chat or show evidence that they are in possession of physical identifying documents (for example, they may ask users to send a selfie holding a matching driver’s license [244]). At times, these methods go too far toward identifying “who” is behind some activity—collecting more information than required to merely verify that there is “a” person conducting the digital activity. Such approaches are not only potentially intrusive from the user’s perspective, but also involve the collection of sensitive personal identifying information, which can introduce important security concerns. Beyond privacy and security issues for users and service providers, such approaches are also not robust to newer AI systems, which are increasingly capable of creating content and avatars that pass these checkpoints. [^39] For an example of a particular case in which even a video call did not successfully head off an AI-powered imitation, see [193]. See also [67, 154]. Beyond using AI to spoof such checks, malicious actors can also purchase photos of people holding up their passports for validation [209].
Finally, digital service providers often use identifiers like emails, phone numbers, and purpose-built hardware-based authenticators to try to verify that there is a person taking actions on their services. For example, one-time passwords sent via SMS can offer some indication that the entity behind some digital activity has access to a physical device—though given the rise of tools for acquiring virtual phone numbers and handling complex workflows through these phone numbers, this indication is not particularly strong. [^40] For instance, Twilio, a company that provides cloud communications services, has been used to acquire phone numbers used in scams [11]. There is a more general problem at play: requiring unique phone numbers, email addresses, and even hardware authenticators at signup can prevent some duplicate account creation, but none of these identifiers are scarce enough to establish that their holder is distinct from other users, and not someone creating deceptive identities at scale. [^41] For more detail on creating fraudulent accounts that require phone number validation, see [248]. Requiring such validation can increase the cost of deception but is not a sufficient solution. Moreover, some of these identifiers facilitate digital tracking [49] and are less private than we aim for with PHCs.
Thus far, we have described trends in AI that could contribute to scalable deception online, and reasons why current solutions may be insufficient for countering it. In the remainder of the paper, we articulate a case for personhood credentials as a foundational tool in our toolkit for protecting a trustworthy digital ecosystem, even as AI grows more advanced.
@@ A PHC system meets the following requirements:
      Issuers check one-per-person requirement at enrollment: The issuer has an effective check of whether a person has already received a personhood credential from them. [^45] Often this check will include some form of evidence that depends on an interaction that occurred in real life, so that AI systems cannot be dispatched to obtain credentials. This evidence need not involve a direct in-person interaction between the user and the issuer. For example, a passport could be used in an enrollment process, providing evidence that its holder at some point went through the necessary in-person steps to acquire such a document. We offer further discussion of methods for satisfying the credential limit in Section C.1.
@@
      Minimal disclosure during usage: When a user presents a personhood credential to a service provider, it reveals to the service provider nothing more than “this person holds a valid PHC” or, with the user’s authorization, “this person holds a valid PHC not yet used with this service.”
@@
      Unlinkability by default: [^49] By unlinkability, we mean unlinkability via direct use of one’s personhood credential. Today, users are sometimes identifiable across websites through factors like their Web browser, even without sharing other identifiers [83]. PHCs are not a direct solution for commonplace tracking of users across the Internet or aggregation of this information via data brokers. However, to the extent that such tracking is facilitated by reusing identifiers across sites (like email addresses or phone numbers), PHCs might make it more difficult to create linkages than previously. By default, service providers or issuers cannot trace or link usage activity across uses, even if issuers and service providers collude. [^50] One risk of government-issued personhood credentials is that the government may be able to compel service providers to turn over records, even if the service provider does not wish to. To be fully privacy-preserving, a personhood credential must be resistant to these forms of subversion. The issuer, by default, learns nothing when a PHC has been used. Service providers do not learn anything when a PHC that has been used with their service is used with another service.

### v2

@@
Anonymity is an important principle online. However, malicious actors have long used misleading identities to conduct fraud, spread disinformation, and carry out other deceptive schemes. With the advent of increasingly capable AI, bad actors can amplify the potential scale and effectiveness of their operations, intensifying the challenge of balancing anonymity and trustworthiness online. In this paper, we analyze the value of a new tool to address this challenge: “personhood credentials” (PHCs), digital credentials that empower users to demonstrate that they are real people—not AIs—to online services, without disclosing any personal information. Such credentials can be issued by a range of trusted institutions—governments or otherwise. A PHC system, according to our definition, could be local or global, and does not need to be biometrics-based. Two trends in AI contribute to the urgency of the challenge: AI’s increasing indistinguishability from people online (i.e., lifelike content and avatars, agentic activity), and AI’s increasing scalability (i.e., cost-effectiveness, accessibility). Drawing on a long history of research into anonymous credentials and “proof-of-personhood” systems, personhood credentials give people a way to signal their trustworthiness on online platforms, and offer service providers new tools for reducing misuse by bad actors. In contrast, existing countermeasures to automated deception—such as CAPTCHAs—are inadequate against sophisticated AI, while stringent identity verification solutions are insufficiently private for many use-cases. After surveying the benefits of personhood credentials, we also examine deployment risks and design challenges. We conclude with actionable next steps for policymakers, technologists, and standards bodies to consider in consultation with the public.
@@
With access to increasingly capable AI, malicious actors can potentially orchestrate more effective deceptive schemes. Two trends contribute to these schemes’ potential impact:
@@
PHCs improve on and complement existing approaches to countering AI-powered deception online. For example, the following approaches are often not robust to highly capable AI, not inclusive, and/or not privacy-preserving:
@@ 1.1   The Internet has long struggled with deceptive activity
Defenders of online freedom of expression sometimes view deception as an unfortunate but necessary cost of preserving users’ privacy [32]. Indeed, the Internet’s pioneers saw anonymity as a fundamental pillar of privacy and freedom of expression [148]. They built its architecture to let people participate in digital spaces without disclosing their real identities. The benefits from this approach are numerous, and worthy of steadfast protection—for example, anonymity allows people in oppressive regimes to express their opinions without fear of retribution. However, these benefits have come at some cost—one such cost is a lack of accountability for deceptive misuse. [^4] For an account of how these norms evolved on Usenet, see [278]. A common approach to undesired content sharing on Usenet was to apply social pressure [21].
@@
The resulting escalation in deceptive activity across social media, public comment systems, and other essential digital services could create substantial challenges for institutions that rely on the Internet. Under pressure, these institutions may resort to invasive measures for verifying users’ identities online, overturning the Internet’s longstanding commitment to privacy and civil liberties. [^9] China’s Ministry of Public Security and its Cyberspace Administration recently proposed a national system to associate online activity with individuals’ legal identities [251]. The proposal has faced criticism from privacy advocates.
@@
In this paper, we focus on one potential solution to counter AI-powered deception: personhood credentials (hereafter referred to as PHCs), which certify that their holder is a person [^10] For the purposes of this paper, we use “person” to refer to a human being, though we recognize that some domains use the term “person” more broadly. For example, in law, “a person is any being whom the law regards as capable of rights or duties” [223], which can include “artificial persons” (e.g., corporations) in addition to “natural persons” (i.e., human beings) [107]. In moral philosophy, “personhood” need not apply only to human beings and could include other groups like non-human animals; “personhood” is sometimes associated with having full moral status, though philosophers disagree considerably about which traits are necessary or sufficient for full moral status [59]. Though we use “person” and “personhood” to refer to human beings, this should not be read as implying any particular view on whether future AI systems should be considered moral persons or granted legal personhood (whether for moral or pragmatic reasons). without revealing anything more about their identity. PHCs help to distinguish people from even the most advanced AI systems by relying on two important deficits of AI. Specifically, AI systems cannot convincingly mimic people offline, and they cannot bypass state-of-the-art cryptographic systems. [^11] Cryptography relies on computationally hard mathematical problems, such as the factoring of very large numbers. There are not any known methods of efficiently solving certain such problems, whether by a human or an AI system. PHCs allow people to interact with different digital service providers in a way that maintains their privacy through unlinkable pseudonymity—the credential holder can have a persistent, privacy-preserving pseudonym with different service providers, and different service providers have no way to link their activity. The issuer of the credential cannot trace their activity either. [^12] While some national digital identity systems also draw on cryptography, they have very different goals and privacy affordances compared to PHCs. For example, Estonia’s national ID-card allows citizens to conduct a wide range of activities online (e.g., cryptographically signing important contracts). But these systems are not intended to conceal which holder of a card is performing the action [213].
@@
That personhood credentials ought to exist is not a wholly novel idea. Related credentials are the subject of ongoing initiatives, [^13] Our advocacy of personhood credentials aligns with many ongoing initiatives that aim for minimal disclosure identity and authentication systems, e.g., World Wide Web Consortium’s (W3C’s) Verifiable Credentials [240] and decentralized identifiers [239], European Union Digital Identity’s (EUDI’s) privacy-preserving digital wallets [88], and a range of standards and implementations of anonymous and attribute-based credentials [1, 126, 90]. One particularly relevant implementation is British Columbia’s Person Credential [34, 165]. as well as decades of research in security and cryptography. In particular, a personhood credential is a type of anonymous credential—first proposed in the 1980s, [^14] An anonymous credential is a credential that allows its holder to prove some specific statement about themselves without revealing anything further [45, 47, 158]. In the case of a personhood credential, the claim is that its holder is a person. In other cases, one could have an anonymous credential that proves their age without revealing anything further. Personhood credentials largely coincide with “accountable pseudonyms,” which aim to link a pseudonymous digital certificate that can be used anonymously online to an offline entity [99]. For a recent systematization of knowledge paper about anonymous credentials, see [118].—and can be understood as similar to a “proof-of-personhood” system. [^15] The term “proof of personhood” emerged in blockchain communities—used in contrast to the concepts of “proof of work” and “proof of stake”—as a strategy for dealing with Sybil attacks (where one entity subverts a digital system by using many pseudonymous identities to exert undue influence) in permissionless voting systems [29]. Given that “proof of personhood systems” often have specific aims (for instance, to allocate a cryptocurrency or voting rights in a decentralized autonomous organization) and different resulting design goals (e.g., they often aim for global uniqueness, sometimes via biometrics), we use the more generic term “personhood credentials” throughout the paper. What is unprecedented is the scale and urgency of the problems these credentials could address, and the harms to trustworthy interactions [^16] By “trustworthy digital interactions,” we mean interactions in which the parties’ reasonable expectations are achieved: for instance, a trustworthy online ecosystem is one in which service providers are confident that their services are being used as intended, and users can access digital services free from fears of abuse, attack, and other harms. We further unpack the foundations of this concept in A What do we mean by “trustworthy” digital interaction?, drawing largely on [133]. on the Internet that may occur if unmitigated.
@@ 1.2   Trends in AI threaten to make online deception more effective
  • Taking human-like actions across the Internet (e.g., browsing websites like an ordinary user, making sophisticated plans to achieve goals they are given, solving CAPTCHAs when challenged).
@@
AI systems are improving in each of these categories. For instance, AI-generated text and images are increasingly difficult to distinguish from human-created content in domains like politics, [^19] For a study of legislators’ response rate to emails drafted by AI compared to those drafted by people—finding only a small difference—see [150]. art [220], and literature [58]. Already, many people struggle to discern whether they are conversing by text with a human or an AI system in some settings. [^20] Research suggests that text written by AI systems are hard to distinguish from text written by people, such as in Turing Test conversations [135, 5]. These findings are not wholly new; at 2020’s GPT-3 launch, OpenAI reported that participants in a research study had trouble distinguishing between news articles by people versus those generated by AI [36]. A growing body of literature explores how AI-generated content might contribute to the efficacy of deceptive schemes, such as through AI’s potential abilities to make persuasive arguments [201, 17, 14, 85], tailor persuasion to a specific recipient [162], and roleplay as certain personas [265].
@@
Likewise, AI is improving at creating human-like avatars. AI can enable malicious actors to create realistic high-quality photos and videos of people who do not exist [147, 33, 259, 35]; animate photos of people into videos [233]; take on the appearance of a different person over video chat [276, 123]; and speak with a realistic human voice, whether a generic voice or that of a specific person. [^21] For an overview of voice generation abilities and resultant challenges, see [196]. There are many different types of risks that can emerge with voice cloning [125]. Voice cloning scams, in particular, are rampant and on the rise [4]. These capabilities can be used to spoof identification checks, such as those requiring a selfie with a matching driver’s license. [^22] Some platform verification processes request proof of a government ID card—but even this process cannot restrict entry of AIs working on behalf of bad actors, as AI tools can be used to spoof a realistic-looking selfie of any person holding their supposed driver’s license [67]. If a service provider is willing to pay for more expensive verifications—such as those offered through the American Association of Motor Vehicle Administrators (AAMVA) in the US—they may be able to filter out some fraudulent IDs, but even these services have limits, such as not offering comparisons against the photo-of-record [7]. In the near future, during a video chat, a user may not be able to discern whether their conversational partner is actually the person they are seeing, a person disguising themselves using AI, or even a complete AI simulation of a real or fictitious person. [^23] There have already been incidents involving convincing deepfaked video calls of executives [50] and politicians [194].
@@ 1.3   Current solutions for countering AI-powered deception need improvement
A second approach is to impose economic barriers that make it costly to perpetuate AI-powered deception at scale. Some websites introduce paid subscriptions with the explicit intent of distinguishing between people and bots by making it more expensive to use many bot accounts. [^35] Elon Musk reported this as one factor in his decision to institute a paid subscription option on Twitter (now X) [224]. These payments, however, disproportionately affect lower-income users and are an insufficient barrier against particularly profitable forms of malicious activity [110, 245]. Relatedly, some digital services check whether users hold a valid credit card—without necessarily charging a payment to the card—though this approach can be circumvented and also has similar challenges with inclusivity. [^36] Credit card requirements can be circumvented via virtual cards [207].
@@
A fourth approach is to use document- or appearance-based verification to confirm that there is a person behind some digital activity. Some verification protocols require that users join a video chat or show evidence that they are in possession of physical identifying documents (for example, they may ask users to send a selfie holding a matching driver’s license [244]). At times, these methods go too far toward identifying “who” is behind some activity—collecting more information than required to merely verify that there is “a” person conducting the digital activity. Such approaches are not only potentially intrusive from the user’s perspective, but also involve the collection of sensitive personal identifying information, which can introduce important security concerns. Beyond privacy and security issues for users and service providers, such approaches are also not robust to newer AI systems, which are increasingly capable of creating content and avatars that pass these checkpoints. [^39] For an example of a particular case in which even a video call did not successfully head off AI-powered deception, see [193]. See also [67, 154]. Beyond using AI to spoof such checks, malicious actors can also purchase photos of people holding up their passports for validation [209].
Finally, digital service providers often use identifiers like emails, phone numbers, and purpose-built hardware-based authenticators to try to verify that there is a person taking actions on their services. For example, one-time passwords sent via SMS can offer some indication that the entity behind some digital activity has access to a physical device—though given the rise of tools for acquiring virtual phone numbers and handling complex workflows through these phone numbers, this indication is not particularly strong. [^40] For instance, Twilio, a company that provides cloud communications services, has been used to acquire phone numbers used in scams [11]. There is a more general problem at play: requiring unique phone numbers, email addresses, and even hardware authenticators at signup can reduce some duplicate account creation, but none of these identifiers are scarce enough to establish that their holder is distinct from other users (and thus that the holder could not be creating deceptive identities at scale). [^41] For more detail on creating fraudulent accounts that require phone number validation, see [248]. Requiring such validation can increase the cost of deception but is not a sufficient solution. Moreover, some of these identifiers facilitate digital tracking [49] and are less private than we aim for with PHCs.
Thus far, we have described trends in AI that could contribute to scalable deception online, as well as reasons why current solutions may be insufficient for countering it. In the remainder of the paper, we articulate a case for personhood credentials as a foundational tool in the toolkit for protecting a trustworthy digital ecosystem, even as AI grows more advanced.
@@ A PHC system meets the following requirements:
      Issuers check one-per-person requirement at enrollment: The issuer has an effective check of whether a person has already received a personhood credential from them. [^45] Often this check will include some form of evidence that depends on an interaction that occurred in real life, so that AI systems cannot be dispatched to obtain credentials. This evidence need not involve a direct in-person interaction between the user and the issuer at the time of enrollment. For example, a passport could be used in an enrollment process, providing evidence that its holder at some point went through the necessary in-person steps to acquire such a document. We offer further discussion of methods for satisfying the credential limit in Section C.1.
@@
      Minimal disclosure during usage: When a user presents evidence of a personhood credential to a service provider, it reveals to the service provider nothing more than “this person holds a valid PHC” or, with the user’s authorization, “this person holds a valid PHC not yet used with this service.”
@@
      Unlinkability by default: [^49] By unlinkability, we mean unlinkability via direct use of one’s personhood credential. Today, users are sometimes identifiable across websites through factors like their Web browser, even without sharing other identifiers [83]. PHCs are not a direct solution for commonplace tracking of users across the Internet or aggregation of this information via data brokers. However, to the extent that such tracking is facilitated by reusing identifiers across sites (like email addresses or phone numbers), PHCs might make it more difficult to create linkages than previously. By default, service providers or issuers cannot trace or link usage activity across uses, even if issuers and service providers collude. [^50] One risk of government-issued personhood credentials is that the government may be able to compel service providers to turn over records, even if a service provider does not wish to. To be fully privacy-preserving, a personhood credential must be resistant to these forms of subversion. The issuer, by default, learns nothing when a PHC has been used. Service providers do not learn anything when a PHC that has been used with their service is used with another service.